OpenStudy (yash2651995):

dipole moment of CHCl3 is less than CH2Cl2 is less than CH3Cl. why?

OpenStudy (anonymous):

CH2Cl2 has the greatest dipole moment. CCl4 is the only one without a dipole. The electronegativity difference in the two atoms causes a polar bond which results in a partial charge on each atom. These molecules are in a tetrahedral configuration, so the two Cl atoms in CH2CL2 do not cancel out, they are additive. These charges act like vectors when in a magnetic field. Think of the geometry or build a model of the molecule and you will see why CHCl3 results in a lower net dipole than CHCl2

OpenStudy (yash2651995):

CH3Cl has maximum dipole (1.87 D) in case of polyatomic molecules(4 atoms one). is how 2 are additive and 3 isnt? all three are at 109.5 degrees from each other one component should add up isnt it?? and CCl4 will be 0 dipole agreed.. resultant of 2 dipoles: Ures=Sqrt(u^2+u^2+u x u x cos (109.5)) =sqrt(u^2(2+cos(109.5)))= u sqrt(2-2*0.325) = u (1.16) which is greater than u ie dipole of one.. so it is additive in case of 2 and even additive in case of three Ures=u(1.694) (warning: maths work could be wrong) then, how can this (what is mentioned in question) happen..

OpenStudy (anonymous):

CH3Cl has greater dipole moment than CH2Cl2 because of symmetry of molecule: